a crane lifts a load upto a height of 8 m in ½ a minute a second crane lifts the same load upto same height in 15 sec .Which crane supplies greater power? Which crane consumes more fuel?

Grade:10

1 Answers

Riddhish Bhalodia
8 years ago
As the height and the mass of the load are same the total work done which is the difference in the energy mgh is the same, how ever
crane1 : t1 = 8/0.5 = 16s
crane2: t2 = 15s

Thus as Power = W/t
crane2 supplies greater power

The fuel consumed is proportional to the work done on the object which is same and hence fuel consumed are also the same